The ratio of the length and width is 4:5 of a rectangle. Find the dimensions if the area is 63 cm squared.
Answer by lwsshak3(11628)

The ratio of the length and width is 4:5 of a rectangle. Find the dimensions if the area is 63 cm squared.
**
let x=length
let y=width
area of a rectangle=length*width
..
x/y=4/5
y=5x/4
..
xy=63
x*5x/4=63
5x^2=252
x^2=252/5
x=√(252/5)
x≈7.1
y≈5x/4≈8.9
length≈7.1 cm
width≈8.9 cm